OpenStudy (wilder.monday):

2.Determine how many, what type, and find the roots for f(x) = x^3 - 5x^2 – 25^x + 125.

OpenStudy (anonymous):

it is a polynomial of degree 3 so there are 3 roots, although two may be complex roots

OpenStudy (anonymous):

you are supposed to find them and you have a lot of fives there, so maybe five is a root

OpenStudy (anonymous):

oh my there is some kind of typo there i think

OpenStudy (wilder.monday):

I tried to like factor it and do all that but I got (x - 5)(x^2 - 25x + 125) and then I don't know how to go on from there

OpenStudy (anonymous):

\[f(x)=x^2+5x^2-25x+125\]

OpenStudy (anonymous):

if you got one root then what remains is a quadratic you can always find the roots of a quadratic, use the formula

OpenStudy (anonymous):

you factored incorrectly is the problem

OpenStudy (anonymous):

\[(x-5)(x^2-25)\] would be better

OpenStudy (wilder.monday):

how did you get that?

OpenStudy (anonymous):

synthetic division is easiest here or you can just think about what it has to be \[x^3+5x^2-25x+125=(x-5)(ax^2+bx+c)\] and \(a, c \) are totally obvious

OpenStudy (anonymous):

since you have to have \(x^3\) then \(a=1\) and since you have to end up with \(125\) you see \(c=-25\) making it \[(x-5)(x^2+bx-25)\] then it will turn out that \(b=0\) giving you \((x-5)(x^2-25)\) or \[(x-5)(x-5)(x+5)\] after you factor the second part

OpenStudy (wilder.monday):

I just did the synthetic division but im not sure what im supposed to get...

OpenStudy (wilder.monday):

English please lol

OpenStudy (anonymous):

you should get as the bottom line 1 0 -25 0

OpenStudy (anonymous):

1 -5 -25 125 5 5 0 -125 _____________________ 1 0 -25 0

OpenStudy (anonymous):

that is what the synthetic division would look like if you are dividing by \(x-5\) the bottom row gives you the quotient of \(x^2-25\)

OpenStudy (wilder.monday):

oooohhh ok I get it. can you help with two more please?

OpenStudy (anonymous):

kk

OpenStudy (wilder.monday):

3.The following graph shows a seventh-degree polynomial: graph of a polynomial that touches the x axis at negative 5, crosses the x axis at negative 1, crosses the y axis at negative 2, crosses the x axis at 4, and crosses the x axis at 7. Part 1: List the polynomial’s zeroes with possible multiplicities. Part 2: Write a possible factored form of the seventh degree function.

OpenStudy (wilder.monday):

this Is only one but theres another question later, but I think its easy

OpenStudy (anonymous):

the zeros you see from your eyeballs right?

OpenStudy (wilder.monday):

the zeros are where the line touches the x-axis right?

OpenStudy (anonymous):

touch or cross yes

OpenStudy (wilder.monday):

how do I make a factored form?

OpenStudy (anonymous):

list them \[{-5,-1,4,7}\] so it will be\[(x+5)^a(x+1)^b(x-4)^c(x-7)^d\] we just need the exponents

OpenStudy (anonymous):

it just touches at \(-5\) and bounced back up, so that exponent is even, no doubt it is 2

OpenStudy (wilder.monday):

the exponents 2?

OpenStudy (anonymous):

at \(-1\) it crosses that exponent is odd, probably 1 now we are at \[(x+5)^2(x+1)(x-4)^c(x-7)^d\]

OpenStudy (anonymous):

the exponent on the factor of \(x+5\) is two yes that is because it touches and bounces back up doesn't cross that is called the "multiplicity" of the zero

OpenStudy (anonymous):

you see it does cross at 7 so that exponent is also 1 \[(x+5)^2(x+1)(x-4)^c(x-7)\]

OpenStudy (wilder.monday):

oooh ok

OpenStudy (anonymous):

and also you see that it crosses at 4, but it is flattened out somewhat there doesn't cross like it does at -5 or at 7 that makes the exponent odd, but not 1, probably 3

OpenStudy (anonymous):

that and the fact that you were told it is a poly of degree 7 we should surmise \[(x+5)^2(x+1)(x-4)^3(x-7)\] which does have degree 7 as asked for

OpenStudy (anonymous):

look at the top picture here to confirm http://www.wolframalpha.com/input/?i=%28x%2B5%29^2%28x%2B1%29%28x-4%29^3%28x-7%29

OpenStudy (wilder.monday):

ok thank you but one question...if the line touches at -5, -1, 4, and 7 how come they are the opposite sign? is it supposed to be like that?

OpenStudy (anonymous):

if you have \[x-3=0\] that meas \[x=3\] right?

OpenStudy (wilder.monday):

oh ok I feel stupid lol I get it now. one more question?

OpenStudy (anonymous):

if you want to solve \[(x-3)(x+4)=0\] that means \[x=3\] or \[x=-4\] conversely if you know the zeros are \(-4\) and \(3\) you know it factors like \[(x-3)(x-4)\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i made a typo on the last post ignore it

OpenStudy (wilder.monday):

4.Without plotting any points other than intercepts, draw a possible graph of the following polynomial: f(x) = (x + 8)^3(x + 6)^2(x + 2)(x – 1)^3(x – 3)^4(x – 6). I know its like the last one and I get how to graph the zeros but I don't know what the line is supposed to lo like.

OpenStudy (wilder.monday):

@satellite73

OpenStudy (anonymous):

crosses at -8 touches but does not cross at -6 crosses at -2 crosses at 1, flattened out a bit touches but does not cross at 3, flattened out a lot crosses at 6

OpenStudy (wilder.monday):

thanks but im not allowed to use a thing that graphs for me. how would I know if I didn't use it?

OpenStudy (anonymous):

by what i wrote above lets take them one at a time

OpenStudy (anonymous):

one factor is \( (x + 8)^3\) the zero is \(-8\) and the multiplicity is 3 3 is odd, so it crosses the axis there but because it is 3 and not one, it is flattened out a bit as it crosses

OpenStudy (anonymous):

next factor is \((x + 6)^2\) the zero is \(-6\) and the multiplicity is 2 because 2 is even, it touches but does not cross there

OpenStudy (anonymous):

etc

OpenStudy (wilder.monday):

ooohhhh ok thanks :)

OpenStudy (kayla2001):

@satellite73 I need help plz